# Super LiDAR Reflectance for Robotic Perception

### v1

@@ IV-B2 Traffic Lane Detection
LiDAR reflectance images reveal a strong contrast between traffic lane markings and the road surface due to differences in physical material and surface properties. By leveraging the dense reflectance images generated by our method, we achieved accurate lane detection under diverse conditions, including very challenging scenarios for cameras (see Fig. 11). The lane detection was performed with LaneATT [49] as the baseline model. The model was trained on the TuSimple dataset [50] and fine-tuned on our dataset, which consists of 1,300 training images, 300 validation images, and 200 test images, with a maximum of four lanes per image. A ResNet-34 backbone [51] was used for fine-tuning the model. The detection performance was evaluated across three distinct environments with varying lighting conditions. The overall detection accuracy ranged from 93.3% to 96.1%, while the false detection rate (FDR) and false negative rate (FNR) remained low, with averages of 6.08% and 6.22%, respectively. The results show the robustness of our approach, which achieves stable and accurate lane detection, particularly under adverse lighting conditions (Fig. 11), outperforming camera-based methods.
@@ V Conclusion
In this work, we introduce a novel benchmark dataset and a framework for generating dense LiDAR reflectance images from sparse measurements, addressing key challenges such as domain adaptation, intensity calibration, and reflectance restoration. Future work will explore generalizing our approach to repeating-scanning LiDAR sensors and balancing real-time performance with densification quality. We also aim to leverage super-resolution reflectance images to enhance robotic perception in dynamic and challenging environments.

### v2

@@ IV-B2 Traffic Lane Detection
LiDAR intensity images reveal a strong contrast between traffic lane markings and the road surface due to differences in surface material physical properties. Leveraging dense intensity images generated by our method, we achieved accurate lane detection under diverse conditions, including challenging scenarios for cameras (Fig. 11). Lane detection was performed with LaneATT [35] as a baseline model. The model was trained on the TuSimple dataset [36] and fine-tuned on our dataset, which consists of 1,300 training images, 300 validation images, and 200 test images, with a maximum of four lanes per image. A ResNet-34 backbone [14] was used for fine-tuning the model. The detection performance was evaluated across three distinct environments with varying lighting conditions. The overall detection accuracy ranged from 93.3% to 96.1%, while the false detection rate (FDR) and false negative rate (FNR) remained low, with averages of 6.08% and 6.22%, respectively. The results show the robustness of our approach, which achieves stable and accurate lane detection, particularly under adverse lighting conditions (Fig. 11), outperforming camera-based methods.
@@ V Conclusion
In this work, we introduce a novel benchmark dataset and a framework for generating dense LiDAR intensity images from sparse measurements, addressing key challenges such as domain adaptation, intensity calibration, and intensity images restoration. Future work will explore generalizing our approach to repeating-scanning LiDAR sensors and balancing real-time performance with densification quality. We also aim to leverage super-resolution LiDAR intensity images to enhance robotic perception in challenging environments.